OpenStudy (anonymous):

find values of a and b such that the function f(x)=ax+b/x^2-1 has a local extreme value of 1 at x=3. is this extreme value a local maxima or minima?

OpenStudy (anonymous):

\[f(x)=\frac{ax+b}{x^2-1}~~?\]

OpenStudy (anonymous):

@nconley or is it \[f(x)=ax+\frac{b}{x^2-1}~~?\] or maybe \[f(x)=ax+\frac{b}{x^2}-1~~?\]